sierra has total of 61 dimes and quarters in her piggybank.
3 more quarters than dimes . the cion have a total value of 10.90 how many dimes and how quters does she have

Answer by addingup(3677)

d+q = 61
q = d+3
0.10d+0.25q = 10.90
---------------------
0.10d+0.25(d+3) = 10.90
0.10d+0.25d+0.75 = 10.90
0.35d = 10.15
d = 29 she has 29 dimes and 61-29 = 32 quarters
-----------------------------
Check:
29(0.10)+32(0.25) = 2.90+8 = 10.90 Correct